OpenStudy (anonymous):

whats the solution for √y+2-6=4

OpenStudy (lilg132):

y + 2 - 6 = 16 16 + 6 - 2 = y 22 - 2 = y 20 = y

OpenStudy (lilg132):

square both sides

OpenStudy (anonymous):

thanks, is that the way i do the problem

OpenStudy (lilg132):

dont forget to click good answer :) and your welcome

OpenStudy (anonymous):

thanks for showing me how to do it than just the answer